40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 901     Entry time: Fri Aug 29 15:01:45 2008
Author: steve 
Type: Update 
Category: PSL 
Subject: MOPA_HTEMP in increasing 
The laser chiller temp is 21.9C ( it should be 20.0C )
Control room temp 73F ok, no obvious block

Ops, there is a piece of paper blocking the intake of the chiller

This is a four day plot. The paper was blocking the air flow all day.
Attachment 1: htcl.jpg  91 kB  | Hide | Hide all
htcl.jpg
ELOG V3.1.3-